# Influence of thermal noise on the field-driven dynamics of the non-collinear antiferromagnet Mn3Sn

### v1

@@ Abstract
In Figs. 3(a) and 3(b), the numerically computed escape time from up state to down state as a function of $H_{\mathrm{ax}}$ and $H_{\mathrm{ay}}$, for five different $\Delta E_{0}$, shown using symbols, is compared to the analytic predictions of Eq. (7) and (8), represented by solid lines. We find that, for both directions of the applied field, the analytical predictions agree very well with the numerical data for $\Delta E_{0}\geq 2k_{\mathrm{B}}T$, though a small discrepancy emerges at higher fields ($H_{\mathrm{a}}\gtrsim 75~{}\mathrm{mT}$) and increases with both $H_{\mathrm{ax}}$ and $H_{\mathrm{ay}}$. This is due to a decrease in the effective barrier heights ($\Delta E_{\mathrm{x,1}}$ and $\Delta E_{y,1}$) below the valid limits of the HTST theory. In the case of $\Delta E_{0}=k_{\mathrm{B}}T$ the discrepancy between the analytic model and numerical data begins at a comparatively lower field ($H_{\mathrm{a}}\gtrsim 45~{}\mathrm{mT}$). For the two directions of magnetic fields, the escape time appears to be more sensitive to $H_{\mathrm{ay}}$ in the low-field regime. This is because an increase in $H_{\mathrm{ay}}$ decreases $\Delta E_{\mathrm{y,1}}$, favoring the octupole transitions from the up state to the down state over both the saddle points equally, thereby speeding up the process. On the other hand, when $H_{\mathrm{ax}}$ increases above zero, thermal field-driven transitions from the up state to the down state are comparatively slower as only $\Delta E_{\mathrm{x,2}}$ decreases but $\Delta E_{\mathrm{x,1}}$ increases. Although not shown here, we find $\tau_{\mathrm{escx,\downarrow}}$ to be same as the results shown in Fig. 3(a), but $\tau_{\mathrm{escy,\downarrow}}$ increases exponentially with $H_{\mathrm{ay}}$ due to increase in $\Delta E_{\mathrm{y,2}}$ (see Supplementary Material).
@@
In conclusion, we comprehensively demonstrated the external field-assisted thermal behaviors of $\mathrm{Mn_{3}Sn}$ within the energy barrier $\leq 5k_{\mathrm{B}}T$. The energy landscape symmetry breakdown is dependent on the field orientation, in which case the theoretical prediction of escape time varies and contributes to two different relaxation modes. The two modes provide prospective applications in random number generation and probabilistic computing while the frequency can reach above GHz. This work focuses on low barrier magnets ($\leq 5k_{\mathrm{B}}T$) and low external field ($\leq 0.1~{}\mathrm{T}$), which is the effective regime of perturbation theory. However, current experiments are reported within high barrier regime ($\geq 40k_{\mathrm{B}}T$) and the switching field is up to $0.5~{}\mathrm{T}$ so our future work will explore the thermal behaviors of $\mathrm{Mn_{3}Sn}$ within this regime. Our work highlights the relevance of thermal noise in establishing the physical understanding and operating principles of scaled $\mathrm{Mn_{3}Sn}$ devices and enable applications such as random number generation and probabilistic computing.
@@ Acknowledgment
Supplementary Material
.1 The Landau-Lifshitz-Gilbert (LLG) equation
The LLG equation for a magnetic moment vector $(\hat{r},\hat{\theta},\hat{\varphi})$ in the spherical coordinates is given as
@@
$H_{\varphi}$ $=-\frac{1}{M_{\mathrm{s}}}\frac{\partial\mathcal{H}}{\partial\varphi},$ (12a)
$H_{\theta}$ $=-\frac{1}{M_{\mathrm{s}}\sin\theta}\frac{\partial\mathcal{H}}{\partial\theta}.$ (12b)
@@
where $\mathcal{M}_{zx}$ is the mirror operation against $zx$ plane and $\mathcal{R}$ is the rotation operation in $xy$ plane, so $\varphi$ in Eq. (11) is replace by $-\varphi$. Owing to the strong exchange interaction between the sublattice vectors, $\mathbf{m}_{i}$ (with $i=1,2,3$), $\mathrm{\theta}_{\mathrm{oct}}$ is restricted around $\pi/2$ and therefore, LLG equation and the effective field for $\mathbf{m}_{\mathrm{oct}}$ is
@@
To average the joint contribution of the three sublattices, the effective field is scaled by a factor of $\frac{1}{3}$e and Liu (2024); Konakanchi et al. (2025) leading to
$H_{\varphi}$ $=-\frac{1}{3M_{\mathrm{s}}}\frac{\partial\mathcal{H}}{\partial\varphi_{\mathrm{oct}}},$ (15a)
$H_{\theta}$ $=-\frac{1}{3M_{\mathrm{s}}\sin\theta_{\mathrm{oct}}}\frac{\partial\mathcal{H}}{\partial\theta_{\mathrm{oct}}}.$ (15b)
@@
$H_{\varphi}$ $=-\frac{1}{3M_{\mathrm{s}}}\frac{\partial\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}}=\frac{1}{2}H_{\mathrm{K}}\sin(2\varphi_{\mathrm{oct}})-\frac{H_{\mathrm{a}}\cos{(\theta_{\mathrm{H}})}}{3(J_{\mathrm{E}}+\sqrt{3}D_{\mathrm{M}})}\Big{[}K_{\mathrm{u}}\sin{(\varphi_{\mathrm{oct}}-\varphi_{\mathrm{H}})}+J_{\mathrm{E}}\delta_{E}\sin{(\varphi_{\mathrm{oct}}+\varphi_{\mathrm{H}})}\Big{]},$ (16a)
$H_{\theta}$ $=-\frac{1}{3M_{\mathrm{s}}}\frac{\partial\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}}=H_{\mathrm{J}}\cos{\theta_{\mathrm{oct}}}\sin{\theta_{\mathrm{oct}}}.$ ) . (16b)
.2 Field-assisted switching
.2.1 $H_{\mathrm{ay}}$-assisted dynamics
@@
With $\varphi_{\mathrm{H}}=3\pi/2$, the Hamiltonian is
in which case local minimum is at $\varphi_{\mathrm{oct}}=\pi/2$ and $23\pi/$ 2, while local maximum is at $\sin\varphi_{\mathrm{oct}}=\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}}$. Using Eq. (17) in Eq. (14), we get
@@
$\dot{\varphi}_{\mathrm{oct}}$ $=\frac{\gamma}{1+\alpha^{2}}\left(-H_{J}\cos\theta_{\mathrm{oct}}\sin\theta_{\mathrm{oct}}+\alpha H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-\alpha H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}}\right).$ (18b)
@@
where $\beta_{y}=3\frac{H_{\mathrm{y,eff}}^{2}}{H_{\mathrm{K}}^{2}}-1$ and the positive eigenvalue of the matrix is
@@
In this work we only analyze $H_{\mathrm{y,eff}}<H_{\mathrm{K}}$ so there is no zero eigenvalue, which indicates no Goldstone mode Rózsa et al. (2019) in this system. So we have $P_{\mathrm{sp}}=P_{\mathrm{min}}=0$ and $V_{\mathrm{min}}/V_{\mathrm{sp}}=1$.
Finally, the escape time with external field applied along $-y$ axis can be evaluated as
@@
.2.2 $H_{\mathrm{ax}}$-assisted dynamics
  • Eigenvalue of linearized LLG
@@
In this case, the LLG in spherical coordinates is given as
@@
$\dot{\varphi}_{\mathrm{oct}}$ $=\frac{\gamma}{1+\alpha^{2}}(-H_{\mathrm{J}}\cos\theta_{\mathrm{oct}}\sin\theta_{\mathrm{oct}}+\alpha H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+\alpha H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}}).$ (25b)
The saddle point corresponding to $\varphi_{\mathrm{oct}}=0$ and $\pi$ give $\cos\varphi_{\mathrm{oct}}=\pm 1$, which implies that the linearized LLG depends on which of the two saddle points are under consideration.
@@
where $\beta_{x}=\mp\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}-1$ for $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively. Therefore, the eigenvalue of the linearized LLG matrix is
@@
Because the saddle points are not identical, the escape processes must consider the two saddle points separately. For the saddle point at $\varphi_{\mathrm{oct}}=0$ and the minimum at $\cos\varphi_{\mathrm{oct}}=-\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}$, the eigenvalues of the Hessian matrix are
@@
$\varepsilon_{\mathrm{1,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=0}=3M_{s}H_{J}\mathcal{V},$ (28b)
@@
$\varepsilon_{\mathrm{2,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=0}=-3M_{s}\mathcal{V}(H_{\mathrm{K}}+H_{\mathrm{x,eff}}),$ (28d)
@@
$\sqrt{\frac{\prod_{j}\left|\varepsilon_{j,\text{sp}}\right|}{\prod_{j}\varepsilon_{j,\min}}}=\sqrt{\quantity|\frac{H_{\mathrm{K}}\quantity(H_{\mathrm{K}}+H_{\mathrm{x,eff}})}{H_{\mathrm{K}}^{2}-H_{\mathrm{x,eff}}^{2}}|}=\sqrt{\quantity|\frac{H_{\mathrm{K}}}{H_{\mathrm{K}}-H_{\mathrm{x,eff}}}|}.$ (29)
Similarly, at the saddle point $\varphi_{\mathrm{oct}}=\pi$ and minimum point $\varphi_{\mathrm{oct}}=-\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}$, the eigenvalues of the Hessian matrix are
$\varepsilon_{\mathrm{1,min}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\cos^{-1}\quantity(\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}})}=3M_{s}H_{J}\mathcal{V},$ (30a)
$\varepsilon_{\mathrm{1,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi}=3M_{s}H_{J}\mathcal{V},$ (30b)
$\varepsilon_{\mathrm{2,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi}=-3M_{s}\mathcal{V}(H_{\mathrm{K}}-H_{\mathrm{x,eff}}),$ (30d)
which gives
$\sqrt{\frac{\prod_{j}\left|\varepsilon_{j,\text{sp}}\right|}{\prod_{j}\varepsilon_{j,\min}}}=\sqrt{\quantity|\frac{H_{\mathrm{K}}(H_{\mathrm{K}}+H_{\mathrm{x,eff}})}{H_{\mathrm{K}}^{2}-H_{\mathrm{x,eff}}^{2}}|}=\sqrt{\quantity|\frac{H_{\mathrm{K}}}{H_{\mathrm{K}}+H_{\mathrm{x,eff}}}|}.$ (31)
@@
.2.3 Analysis of depopulation factor

### v2

@@ Abstract
In Figs. 3(a) and 3(b), the numerically computed escape time from up state to down state as a function of $H_{\mathrm{ax}}$ and $H_{\mathrm{ay}}$, for five different $\Delta E_{0}$, shown using symbols, is compared to the analytic predictions of Eq. (7) and (8), represented by solid lines. We find that, for both directions of the applied field, the analytical predictions agree very well with the numerical data for $\Delta E_{0}\geq{3}k_{\mathrm{B}}T$, though a small discrepancy emerges at higher fields ($H_{\mathrm{a}}\gtrsim 75~\mathrm{mT}$) and increases with both $H_{\mathrm{ax}}$ and $H_{\mathrm{ay}}$. This discrepancy arises from a reduction in the effective barrier heights ($\Delta E_{\mathrm{x,1}}$ and $\Delta E_{\mathrm{y,1}}$) below the valid limits of HTST. In addition, increasing the field may further amplify the deviation of Eq. (2) from Eq. (1), ultimately leading to inaccuracies in $\tau_{\mathrm{esc}}$. A detailed discussion of this issue is provided in Sec. I.3.1 of the Supplementary Material.
For the two directions of magnetic fields, the escape time appears to be more sensitive to $H_{\mathrm{ay}}$ in the low-field regime. This is because an increase in $H_{\mathrm{ay}}$ decreases $\Delta E_{\mathrm{y,1}}$, favoring the octupole transitions from the up state to the down state over both the saddle points equally, thereby speeding up the process. On the other hand, when $H_{\mathrm{ax}}$ increases above zero, thermal field-driven transitions from the up state to the down state are comparatively slower as only $\Delta E_{\mathrm{x,2}}$ decreases but $\Delta E_{\mathrm{x,1}}$ increases. The sensitivity of escape time to various magnetic parameters is quantified in Sec. I.3.2 of the Supplementary Material. Although not shown here, we find $\tau_{\mathrm{escx,\downarrow}}$ to be same as the results shown in Fig. 3(a), but $\tau_{\mathrm{escy,\downarrow}}$ increases exponentially with $H_{\mathrm{ay}}$ due to increase in $\Delta E_{\mathrm{y,2}}$ as shown in Sec. I.3.3 of the Supplementary Material.
@@
In conclusion, we comprehensively demonstrated the external field-assisted thermal behaviors of $\mathrm{Mn_{3}Sn}$ for energy barriers in the range of $(3-6)k_{\mathrm{B}}T$. The energy landscape symmetry breakdown is dependent on the field orientation, in which case the theoretical prediction of escape time varies and contributes to two different relaxation modes. The two modes provide prospective applications in random number generation and probabilistic computing while the frequency can reach above GHz. This work focuses on low barrier magnets $\left((3-6)k_{\mathrm{B}}T\right)$ and low external field ($\leq 0.1~\mathrm{T}$), which is the effective regime of perturbation theory. However, current experiments are reported for high barrier regime ($\geq 40k_{\mathrm{B}}T$) and the switching field is up to $0.5~\mathrm{T}$ so our future work will explore the thermal behaviors of $\mathrm{Mn_{3}Sn}$ within this regime. Our work highlights the relevance of thermal noise in establishing the physical understanding and operating principles of scaled $\mathrm{Mn_{3}Sn}$ devices and enable applications such as random number generation and probabilistic computing.
@@ Acknowledgment
I Supplementary Material
I.1 Analytical derivation of escape time
Following steps must be implemented to derive Eq. (5) in the main text.:
  • Linearizing the equations describing the motion of $\mathbf{m}_{\mathrm{oct}}$ to get the only positive eigenvalue $\lambda_{+}$ at the saddle point.
  • Evaluating the Hessian matrix to obtain the eigenvalues $\varepsilon_{\mathrm{j,min}}$ at the energy minimum and $\varepsilon_{\mathrm{j,sp}}$ at the saddle point. The number of eigenvalues equal to zero determines the number of Goldstone states and the associated parameters such as $V_{\mathrm{min}}$, $P_{\mathrm{min}}$, $V_{\mathrm{sp}}$, and $P_{\mathrm{sp}}$.
  • Calculating the depopulation factor $A(\frac{\delta E}{k_{\mathrm{B}}T})$ through numerical analysis.
I.1.1 The Landau-Lifshitz-Gilbert (LLG) equation
In preparation for the linearization of the LLG equation describing the motion of $\mathbf{m}_{\mathrm{oct}}$, we first derive the equations directly in terms of $\mathbf{m}_{\mathrm{oct}}=(\sin\theta_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}},\sin\theta_{\mathrm{oct}}\sin\varphi_{\mathrm{oct}},\cos\theta_{\mathrm{oct}})$ where $\varphi_{\mathrm{oct}}$ and $\theta_{\mathrm{oct}}$ are the azimuthal and polar angles of the octupole moment as depicted in Fig. 1(b) in the main text; therefore the equations of motion are explicitly dependent on $\theta_{\mathrm{oct}}$ and $\varphi_{\mathrm{oct}}$.
[…]
@@ I.1.1 The Landau-Lifshitz-Gilbert (LLG) equation
$H_{\varphi}$ $=-\frac{1}{M_{\mathrm{s}}\sin\theta}\frac{\partial\mathcal{H}}{\partial\varphi},$ (12a)
$H_{\theta}$ $=-\frac{1}{M_{\mathrm{s}}}\frac{\partial\mathcal{H}}{\partial\theta}.$ (12b)
@@
where $\mathcal{M}_{zx}$ is the mirror operation against $zx$ plane and $\mathcal{R}$ is the rotation operation in $xy$ plane, so $\varphi$ in Eq. (11) is replace by $\varphi$. Owing to the strong exchange interaction between the sublattice vectors, $\mathbf{m}_{i}$ (with $i=1,2,3$), $\mathrm{\theta}_{\mathrm{oct}}$ is restricted around $\pi/2$ and, therefore, the LLG equation and the effective field for $\mathbf{m}_{\mathrm{oct}}$ is
@@
To average the joint contribution of the three sublattices, the effective field is scaled by a factor of $\frac{1}{3}$He and Liu (2024); Konakanchi et al. (2025) leading to
$H_{\varphi}$ $=-\frac{1}{3M_{\mathrm{s}}\sin\theta_{\mathrm{oct}}}\frac{\partial\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}},$ (15a)
$H_{\theta}$ $=-\frac{1}{3M_{\mathrm{s}}}\frac{\partial\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}}.$ (15b)
@@
[…]
I.1.2 Escape time with $H_{\mathrm{ay}}$
With $\varphi_{\mathrm{H}}=3\pi/2$, the Hamiltonian from Eq. (2) in the main text is
in which case local minimum is at $\varphi_{\mathrm{oct}}=\pi/2$ and $3\pi/2$, while local maximum is at $\sin\varphi_{\mathrm{oct}}=\frac{H_{\mathrm{y,eff}}}{H_{\mathrm{K}}}$.
@@ I.1.2 Escape time with $H_{\mathrm{ay}}$
Substituting Eq. (17) into Eq. (14) and Eq. (16), the LLG equation of $\mathbf{m}_{\mathrm{oct}}$ is written as
@@
$\dot{\varphi}_{\mathrm{oct}}$ $=\frac{\gamma}{{(}1+\alpha^{2}{)\sin\theta_{\mathrm{oct}}}}\left(-H_{J}\cos\theta_{\mathrm{oct}}\sin\theta_{\mathrm{oct}}+\alpha H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}-\alpha H_{\mathrm{y,eff}}\cos\varphi_{\mathrm{oct}}\right).$ (18b)
@@
where $\beta_{y}=3\frac{H_{\mathrm{y,eff}}^{2}}{H_{\mathrm{K}}^{2}}-1$. The positive eigenvalue of the matrix is
@@
In this work, we only analyze $H_{\mathrm{y,eff}}<H_{\mathrm{K}}$ so there is no zero eigenvalue, which indicates no Goldstone mode Rózsa et al. (2019) in this system. So we have $P_{\mathrm{sp}}=P_{\mathrm{min}}=0$ and $V_{\mathrm{min}}/V_{\mathrm{sp}}=1$.
Finally, the escape time with external field applied along y-y axis can be evaluated as
@@
I.1.3 Escape time with $H_{\mathrm{ax}}$
@@
  • Eigenvalue of linearized LLG
Substituting Eq. 23 into Eq. (14) and Eq. (16),the LLG equations of $\mathbf{m}_{\mathrm{oct}}$ is written as:
@@
$\dot{\varphi}_{\mathrm{oct}}$ $=\frac{\gamma}{{(}1+\alpha^{2}{)\sin\theta_{\mathrm{oct}}}}(-H_{\mathrm{J}}\cos\theta_{\mathrm{oct}}\sin\theta_{\mathrm{oct}}+\alpha H_{\mathrm{K}}\sin\varphi_{\mathrm{oct}}\cos\varphi_{\mathrm{oct}}+\alpha H_{\mathrm{x,eff}}\sin\varphi_{\mathrm{oct}}).$ (25b)
Unlike the case with $H_{\mathrm{ay}}$, the two saddle points are not degenerate. Specifically, $\cos\varphi_{\mathrm{oct}}=1$ and 1-1 correspond to two distinct saddle points at $\varphi_{\mathrm{oct}}=0$ and $\pi$, respectively. As a result, the corresponding matrices and their eigenvalues must be evaluated separately for each saddle point.
@@
where $\beta_{x}$ is $\varphi_{\mathrm{oct}}$ dependent:
$\beta_{x}=\begin{cases}-\frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}-1\ \ \ &\varphi_{\mathrm{oct}}=0\\[10.0pt] \frac{H_{\mathrm{x,eff}}}{H_{\mathrm{K}}}-1\ \ \ \ &\varphi_{\mathrm{oct}}=\pi\end{cases}$ (28)
@@
Because the saddle points are not identical, the eigenvalues of Hessian matrix with $H_{\mathrm{ax}}$ need to be evaluated separately:
@@
$\varepsilon_{\mathrm{1,sp}}$ $=\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\theta_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=0{/\pi}}=3M_{s}H_{J}\mathcal{V},$ (29b)
@@
${\varepsilon_{\mathrm{2,sp}}}$ ${=}\begin{cases}{\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=0}=-3M_{s}\mathcal{V}(H_{\mathrm{K}}+H_{\mathrm{x,eff}})\ \ \ \varphi_{\mathrm{oct}}=0,}\\ {\mathcal{V}\frac{\partial^{2}\mathcal{H}_{\mathrm{oct}}}{\partial\varphi_{\mathrm{oct}}^{2}}\Big{|}_{\theta_{\mathrm{oct}}=\pi/2,\varphi_{\mathrm{oct}}=\pi}=-3M_{s}\mathcal{V}(H_{\mathrm{K}}-H_{\mathrm{x,eff}})\ \ \ \varphi_{\mathrm{oct}}=\pi,}\end{cases}$ (29d)
@@
$\sqrt{\frac{\prod_{j}\left|\varepsilon_{j,\text{sp}}\right|}{\prod_{j}\varepsilon_{j,\min}}}=\begin{cases}\sqrt{\quantity|\frac{H_{\mathrm{K}}\quantity(H_{\mathrm{K}}+H_{\mathrm{x,eff}})}{H_{\mathrm{K}}^{2}-H_{\mathrm{x,eff}}^{2}}|}=\sqrt{\quantity|\frac{H_{\mathrm{K}}}{H_{\mathrm{K}}-H_{\mathrm{x,eff}}}|},\\[20.0pt] \sqrt{\quantity|\frac{H_{\mathrm{K}}(H_{\mathrm{K}}+H_{\mathrm{x,eff}})}{H_{\mathrm{K}}^{2}-H_{\mathrm{x,eff}}^{2}}|}=\sqrt{\quantity|\frac{H_{\mathrm{K}}}{H_{\mathrm{K}}+H_{\mathrm{x,eff}}}|}\end{cases}$ (30)
@@
I.1.4 Analysis of depopulation factor